# OpenGait: A Comprehensive Benchmark Study for Gait Recognition towards Better Practicality

### v1

@@ II-D Other Related Works
Survey on Gait Recognition. While survey papers [39, 40, 22] have offered comprehensive overviews of gait recognition publications, a detailed experimental analysis is still lacking. In contrast, several impactful works in other fields, such as recommendation system [41], metric learning [42], and unsupervised domain adaptation [43], have revisited SoTA experimentally. Recognizing this gap, OpenGait aims to fill this void by not only re-evaluating recent works but also uncovering new insights in the field of gait recognition.
Codebase for Computer Vision. In the computer vision community, having a robust codebase is crucial for advancing research in specific domains. For example, Amos et al. [44] proposed OpenFace, a face recognition library that bridges the gap between public face recognition systems and industry-leading private systems. In the field of object detection, a PyTorch toolbox called MMDetection [45] supports almost all popular detection methods, providing a convenient platform for systematic comparison. As gait recognition continues to evolve rapidly, the need for an infrastructure code platform has become increasingly evident.
III A Glance of OpenGait Platform
Over the past few years, numerous new methods and datasets have emerged for gait recognition. However, the lack of a unified and fair evaluation platform cannot be overlooked. To facilitate academic research, this paper presents a PyTorch-based [46] toolbox, termed OpenGait, as a reasonable and dependable solution to address this issue.
@@ III-A Design Principles of OpenGait
As shown in Fig. 2, OpenGait covers the highlights below.
Compatibility with Diverse Gait Modalities. Typical gait modalities include the silhouette image, 2D/3D skeleton, and emerging ones like the SMPL model [27, 9], point cloud [21], and RGB image [24, 26]. While existing open-source repositories mostly support one specific input type, OpenGait is designed to support all of these representations.
Compatibility with Various Frameworks. The landscape of gait recognition is evolving with the emergence of novel frameworks, such as multi-modal [9], end-to-end [24, 27, 26], and contrastive learning [37] based methods. Unlike existing repositories that are specifically tailored to certain models, OpenGait is designed to be flexible and extensible, accommodating all of the above frameworks seamlessly.
Support for Various Evaluation Datasets. OpenGait offers compatibility with commonly used gait datasets, covering a wide range of scenarios and sensors. It seamlessly integrates with constrained datasets like CASIA-B [10] and OU-MVLP [11], as well as in-the-wild datasets GREW [8] and Gait3D [9]. Moreover, OpenGait also supports the latest in-the-lab gait datasets such as CCPG [20] and SUSTech1K [21]. Notably, OpenGait’s compatibility has been extended to major competitions like HID [13] and GREW [8], where it has been instrumental in the development of winning solutions.
Support for SoTA. OpenGait encompasses the reproduction of most SoTA methods mentioned in the related work section. The reproduced performances match or surpass the results reported by the original papers. These rich official examples help beginners get started with the code better and faster. Moreover, this also provides a solid platform for fair and thorough comparisons in this benchmark study.
III-B Main Modules
We follow the design of most PyTorch deep learning projects and divide OpenGait into three modules, i.e., data, modeling, and evaluation, as shown in Fig. 2.
Data module contains the data loader, sampler, and transform scripts, which are responsible for loading, sampling, and pre-processing the input data flow respectively.
@@ III-B Main Modules
IV State-of-the-art Methods Revisiting
With the help of OpenGait, we conduct a comprehensive re-evaluation of several SoTA gait methods. Through meticulous ablation studies, we uncover some interesting insights that differ from those presented in the original papers.
@@ IV-A Experimental Recheck on Representative Methods
A lot of prior works only perform experiments on indoor gait datasets, notably CASIA-B [10] and OU-MVLP [11], with further ablation studies typically limited to CASIA-B [10]. This subsection aims to extend the analysis by conducting additional ablation studies on an in-the-wild dataset, Gait3D [9], to assess their robustness to real-world gait data.
TABLE II: The effect of MGP and Multi-scale HPP in GaitSet [7].
@@
Re-conduct Ablation Study on GaitSet. With taking the silhouettes as input, GaitSet [7] treats the gait sequence as an unordered set and uses a simple maximum pooling function along the temporal dimension, called Set Pooling (SP), to generate the set-level understanding of the entire input video. GaitSet [7] provides insights for many subsequent works thanks to its simplicity and effectiveness. However, we find that the other two important components in GaitSet [7], namely the parallel Multi-layer Global Pipeline (MGP) and pyramid-like Horizontal Pyramid Pooling (HPP) [47], do not work well enough on both the indoor CASIA-B and outdoor Gait3D datasets. Specifically, as shown in Fig. 3 (a), MGP can be regarded as an additional branch that aggregates the hierarchical set-level characteristics. HPP [47] follows the fashion feature pyramid structure aiming to extract multi-scale part-based features. As shown in Table II, if we strip MGP or remove the multi-scale mechanism in HPP from the official GaitSet [7], the obtained model could reach the same or even better performance on both CASIA-B and Gait3D with saving over 80% training weights. This result indicates that the set-level characteristics extracted by bottom convolution blocks may be hard to benefit the final gait representation666 In the original GaitSet [7], all the obtained partial vectors will be concatenated into single feature vector used for final evaluation. However, the followup works [15, 16, 9] find this concatenation unnecessary and take the average of partial distance as a final metric. Therefore, OpenGait follows this manner and reproduces a better performance than official results. . Besides, the multi-scale mechanism in HPP also provides no extra discriminative features. The cause may be that the employed statistical pooling functions are too weak to learn extra knowledge from various-scale human body parts.
Figure 3: The key modules of four previous SoTA methods. The modules enclosed by dotted green lines are the ones we remove or replace for comparison. (a): MGP and multi-scale mechanism in HPP are removed [7]. (b): FConvs are replaced with the regular convolution layers [15]. (c): The upper local feature branch is replaced with an independent conv3d that is identical to the lower global branch [16]. (d): SLN branch is replaced with our stronger backbone [17].
@@
Re-conduct Ablation Study on GaitPart. One of the core contributions of GaitPart [15] is to point out the importance of local details with the proposed Focal Convolution (FConv) layer. Fig. 3 (b) shows the receptive field’s expansion of the top-layer neuron in the network established by regular vs. focal convolution layers. Technically, FConv splits the input feature map into several parts horizontally and then performs a regular convolution over each part separately. As shown in Table III, we get a much higher performance (Rank-1: +7.0%) on Gait3D by degenerating the FConv to the regular convolution layer. This phenomenon exhibits that the extraction of gait features may be seriously affected by merely splitting the feature map in a fixed-window manner, due to the low-quality segmentation of wild data. The shifted window mechanism [48] may offer a solution to this issue.
@@
Re-conduct Ablation Study on GaitGL. As shown in Fig 3 (c), GaitGL [16] develops the global and local convolution layer, where the local branch can be regarded as the FConv [15] employing 3D convolution while the global branch presents a standard 3D convolution layer. Similar to GaitPart [15], as shown in Table IV, removing the local branch can achieve better performance on the Gait3D dataset.
Re-conduct Ablation Study on SMPLGait. As shown in Fig. 3 (d), SMPLGait [9] consists of two elaborately-designed branches, i.e., silhouette (SLN) and SMPL (3D-STN) branch. They are respectively used for 2D appearance extraction and 3D knowledge learning. SMPLGait [9] takes advantage of the 3D mesh data available in Gait3D [9] and achieves performance gains on the top of the silhouette branch. However, Table V demonstrates that the proposed SMPL branch does not provide obvious benefits when we give the silhouette branch a strong backbone network like GaitBase [17].
@@
SLN and ResNet9 respectively denote the backbone used by SMPLGait [9] and GaitBase [17].
In our view, there are three potential reasons causing the failure of the SMPL branch: a) Though the SMPL model is usually visualized as a dense mesh, its feature vector only possesses tens of dimensions that present the relatively sparse characterization of body shape and posture, making it challenging to enhance the fine-grained description of gait patterns. b) Since the SMPL model is not recognition-oriented, purposefully fine-tuning it may be more optimal than directly utilizing it to depict the subtle individual characteristics [27]. c) In the wild, estimating an accurate SMPL model that finely captures body shape and posture from a single RGB camera is still challenging. In a nutshell, introducing 3D geometrical information from the SMPL model to enhance gait representation learning is well worth further exploration.
@@ IV-B Rethinking and Discussion
Drawing from the aforementioned findings, we propose that the model structures of previous gait recognition methods may lack robustness. This limitation stems, largely, from the heavy reliance on constrained datasets that struggled with simulating the complexity of outdoor environments. To provide a comprehensive understanding, we elaborate on this point below.
@@
Necessity of Outdoor Evaluation. Previous methods are primarily evaluated on the indoor CASIA-B [10] and OU-MVLP [11]. We argue that this practice suffers from three significant drawbacks: a) Indoor settings. The walking videos are captured by a camera array and subjects are requested to follow a particular route. It makes the data source significantly different from real-world scenarios. b) Simple background. The simple laboratory background cannot reflect the complex changes of wild scenes. c) Outdated processing methods. The RGB videos are processed by the background subtraction algorithm significantly distinct from the deep-learning-based segmentation algorithms used for current applications.
Recently, there has been a noticeable emergence of large-scale real-world gait datasets [8, 9] aimed at advancing gait recognition from controlled laboratory settings to real scenarios. However, despite these advancements, many of the lately-published works [49, 50, 51] continue to heavily rely on constrained setups. To advance the development of gait recognition systems, we advocate for increased attention and utilization of outdoor gait datasets.
Necessity of Sufficient Ablative Experiments. The ablation study is recognized as the primary means of evaluating the effectiveness of individual components within a proposed method. However, we consider that the conclusions drawn from ablative experiments performed merely on CASIA-B [10] may lead to limited practical applicability. The primary reasons are two-fold: a)  CASIA-B contains only 50 subjects for evaluation. Such few testing participants make the results vulnerable to noisy factors. b)  In contemporary applications, pedestrian segmentation predominantly relies on deep learning models, diverging from the outdated background subtraction algorithm employed by CASIA-B. In principle, performing a comprehensive ablation study across diverse large-scale datasets can yield a more robust configuration of hyperparameters and the design of model architectures.
Necessity of A Strong Backbone. The quality of a model largely hinges on the capability of its backbone, and a sub-optimal backbone may inflate the effectiveness of additional modules. The evolution of CNN architectures has progressed from shallow to deep, giving rise to the emergence of excellent backbone networks such as AlexNet [52], VGG-16 [53], and ResNet [18]. However, previous works [7, 54, 9] in gait recognition have predominantly relied on plain convolutional neural networks, composed of several basic convolution layers. As gait recognition research progresses towards applications and large-scale real-world datasets [8, 9] become more accessible, the necessity for a robust and powerful backbone network becomes apparent and essential for achieving accurate and dependable gait recognition results.
@@
Figure 4: (a) Overall architecture. The TP and HP respectively denote Temporal and Horizontal Pooling. The Head comprises the separate fully-connected and BNNeck [55] layers. (b-d) Employed basic blocks, including the (b) 2D, (c) 3D and (d) Pseudo 3D residual convolution blocks [18, 56, 57].
V Gait Recognition Baselines Reconstruction
@@
Building on the insights outlined earlier, this section establishes a series of baseline models aimed at uncovering the practical principles for contemporary in-the-wild gait recognition. Our investigation encompasses appearance-based, model-based, and multi-modal methodologies, enriching the comprehensiveness of this benchmark study. To ensure broader applicability, we prioritize the development of structurally simple yet experimentally powerful and empirically robust architectures to lay the foundations for follow-ups. To this end, we undertake three-fold efforts:
@@
  • A General Pipeline. To ensure the baseline models’ typicality and universality, we adopt a widely recognized pipeline as shown in Fig. 4 (a).
@@
  • Two Basic Blocks. To make it simple and scaleable, we employ only two blocks illustrated in Fig. 4 (b), i.e., the 3D residual unit and its pseudo counterpart [56, 57] to build the network backbone.
@@
  • Three Strong Baselines. As depicted in Fig. 4 (c), the resulting DeepGaitV2, SkeletonGait, and SkeletonGait++ share an identical pipeline, each respectively taking the silhouette image, skeletal coordinates, and their combination as input, accordingly representing the appearance-based, model-based, and multi-modal baseline models in this benchmark study. Notably, we align the data format of silhouette and skeleton by drawing the latter’s coordinates into an imagery heatmap, facilitating flexible comparison and fusion. Additional details will be provided in Sec. V-C.
The subsequent content of this section begins by introducing the used pipeline. Following this, we present DeepGaitV2 based on a large-scale experimental investigation on both in-the-lab and in-the-wild gait datasets. For brevity, implementation details are temporarily omitted, and they will be comprehensively presented in Sec. VI. We then proceed to introduce SkeletonGait and SkeletonGait++ in detail.
V-A Pipeline

### v2

@@ II-D Other Related Works
Survey on Gait Recognition. While survey papers [46, 47, 24] have offered comprehensive overviews on gait recognition, detailed experimental analysis and comparisons are still lacking. In contrast, several impactful works in other fields, such as recommendation system [48], metric learning [49], and unsupervised domain adaptation [50], have revisited SoTA experimentally. Similarly, OpenGait aims to do this by not only re-evaluating recent methods but also uncovering some new insights in gait recognition.
Codebase for Computer Vision. In the computer vision community, a robust codebase is crucial for advancing research in specific domains. For instance, Amos et al. [51] proposes OpenFace, a face recognition library that bridges the gap between public face recognition systems and industry-leading private systems. In object detection, mmdetection [52] supports almost all popular detection methods, providing a convenient platform for systematic comparison. As gait recognition continues to evolve rapidly, the need for an infrastructure code platform has become increasingly evident.
III A Glance at OpenGait Platform
Over the past few years, numerous new methods and datasets have emerged for gait recognition. However, the lack of a unified and fair evaluation platform cannot be overlooked. A PyTorch-based [53] toolbox, termed OpenGait, is presented in the following as a solution to this issue.
@@ III-A Design Principles of OpenGait
We have considered the following design principles below when we designed OpenGait.
Compatibility with Multiple Gait Modalities. Typical gait modalities include silhouette images, 2D/3D skeletons, and emerging ones like the SMPL model [32, 9], point clouds [21], and RGB images [29, 31]. Most open-source projects are designed for specific methods and mostly supporting just one or a limited number of gait modalities. In this work, OpenGait tries to support all popular modalities.
Compatibility with Various Frameworks. The landscape of gait recognition is evolving with the emergence of novel frameworks, such as multi-modal [9], end-to-end [29, 32, 31], and contrastive learning [44] based methods. Unlike some open-source projects specifically tailored to certain models, OpenGait is designed to be flexible and extensible, and can accommodate all of the above frameworks seamlessly.
Support for Various Evaluation Datasets. OpenGait offers compatibility with commonly used gait datasets, which cover a wide range of scenarios and sensors. It seamlessly integrates with indoor datasets like CASIA-B [10] and OU-MVLP [11], as well as in-the-wild datasets, such as GREW [8] and Gait3D [9]. Moreover, OpenGait also supports the latest in-the-lab gait datasets such as CCPG [20] and SUSTech1K [21]. Thanks to its flexibility and reusability, OpenGait has been extended to major gait recognition competitions like HID [13] and GREW [8], where it has been instrumental in the development of winners’ solutions.
Support for SoTA. OpenGait encompasses the reproduction of most SoTA methods mentioned and will keep updated to date by appending promising methods. The results by the reproduced methods match or surpass the ones reported in the original papers. In this way, we provided rich examples to help beginners get started with the code. Moreover, OpenGait also presents a solid platform for fair and thorough comparisons in this benchmark study.
III-B Modules in OpenGait
We follow the designs of a typical PyTorch deep learning project and divide OpenGait into three modules, i.e., data, modeling, and evaluation, as shown in Fig. 2.
Data module contains some data loaders, data samplers, and data transform scripts, and they are for loading, sampling, and pre-processing the input data, respectively.
@@ III-B Modules in OpenGait
OpenGait was developed in Python language, and easy to use. Users can easily train a new deep model on multiple datasets by modifying some configurations. Its GitHub repository [^6] https://github.com/ShiqiYu/OpenGait has gained 659 stars and 153 forks up to June 2024.
IV Revisiting of State-of-the-art Methods
With the help of OpenGait, we can conduct a comprehensive re-evaluation of several representative gait methods efficiently. Through meticulous ablation studies, some interesting insights, which are different from the ones presented in the original papers, have been recovered.
@@ IV-A Experimental Recheck on Representative Methods
A lot of works only perform experiments on indoor gait datasets, notably CASIA-B [10] and OU-MVLP [11], with further ablation studies typically limited on CASIA-B [10]. This subsection aims to extend some additional ablation studies to a large in-the-wild dataset, Gait3D [9], to assess the methods’ robustness in real-world scenarios.
TABLE II: The effectiveness of MGP and Multi-scale HPP in GaitSet [7].
@@
Re-conduct Ablation Study of GaitSet. With taking the silhouettes as input, GaitSet [7] treats the gait sequence as an unordered set and uses a simple maximum pooling function along the temporal dimension, called Set Pooling (SP), to extract a set-level understanding of the entire input video. GaitSet [7] provides insights for many subsequent works for its simplicity and effectiveness. However, we find that two important components in GaitSet [7], namely the parallel Multi-layer Global Pipeline (MGP) and pyramid-like Horizontal Pyramid Pooling (HPP) [54], do not work well enough on the indoor CASIA-B and outdoor Gait3D datasets. Specifically, as shown in Fig. 3 (a), MGP can be regarded as an additional branch that aggregates the hierarchical set-level characteristics. HPP [54] follows the fashion of feature pyramid structure, and tries to extract multi-scale part-based features. As shown in Table II, if we strip the MGP or remove the multi-scale mechanism in HPP from the original GaitSet [7], the revised model can reach the same or even better performance on both CASIA-B and Gait3D with saving over 80% training weights. This result indicates that the set-level characteristics extracted by the bottom convolution blocks may be difficult to benefit the final gait representation777 In the original GaitSet [7], all the obtained partial vectors will be concatenated into a single feature vector used for final evaluation. However, the follow-up works [15, 16, 9] find this concatenation unnecessary and take the partial distance as a final metric. Therefore, OpenGait follows this manner and reproduces a better performance than the original GaitSet. . Besides, the multi-scale mechanism in HPP cannot provide discriminative features. The cause may be that the employed statistical pooling functions are too weak to learn extra knowledge from various-scale human body parts.
Figure 3: The key modules of four SoTA methods. The modules enclosed by dotted green lines are the ones removed or replaced for comparisons. (a): MGP and multi-scale mechanism in HPP are to be removed [7]. (b): FConvs are to be replaced with the regular convolution layers [15]. (c): The upper local feature branch is to be replaced with an independent conv3d that is identical to the lower global branch [16]. (d): SLN branch is to be replaced with our strong backbone [17].
@@
Re-conduct Ablation Study of GaitPart. One of the core contributions of GaitPart [15] is to point out the importance of local details with the proposed Focal Convolution (FConv) layer. Fig. 3 (b) shows the receptive field’s expansion of the top-layer neuron in the network established by regular vs. focal convolution layers. Technically, FConv splits the input feature map into several parts horizontally and then performs a regular convolution over each part separately. As shown in Table III, we get much better performance (Rank-1: +7.0%) on Gait3D by degenerating FConv into the regular convolution layer. This phenomenon exhibits that the extraction of gait features may be seriously affected by merely splitting the feature map in a fixed-window manner, due to the low-quality segmentation of the outdoor data. The shifted window mechanism [55] may offer a better solution to this issue.
@@
Re-conduct Ablation Study of GaitGL. As shown in Fig 3 (c), GaitGL [16] develops the global and local convolution layer, where the local branch can be regarded as the FConv [15] employing 3D convolution while the global branch presents a standard 3D convolution layer. Similar to GaitPart [15], as shown in Table IV, removing the local branch can achieve better performance on the Gait3D dataset.
Re-conduct Ablation Study of SMPLGait. As shown in Fig. 3 (d), SMPLGait [9] consists of two elaborately-designed branches, i.e., the silhouette (SLN) and SMPL (3D-STN) branch, respectively using for 2D appearance extraction and 3D knowledge learning. SMPLGait [9] takes advantage of the 3D mesh data available in Gait3D [9] and achieves performance gains on the top of the silhouette branch. However, Table V demonstrates that the proposed SMPL branch does not provide obvious benefits when we give the silhouette branch a strong backbone network like GaitBase [17].
@@
Note: SLN and ResNet9 denote the backbone used by SMPLGait [9] and GaitBase [17], respectively.
In our view, there should be three potential reasons causing the failure of the SMPL branch: a) Though the SMPL model is usually visualized as a dense mesh, its feature vector only possesses tens of dimensions that present the relatively sparse characterization of body shape and posture, making it challenging to enhance the fine-grained description of gait patterns. b) Since the SMPL model is not recognition-oriented, purposefully fine-tuning it may be more optimal than directly utilizing it to depict the subtle individual characteristics as in [32]. c) In the wild, estimating an accurate SMPL model that finely captures body shape and posture from a single RGB camera is still challenging. In a nutshell, introducing 3D geometrical information from the SMPL model to enhance gait representation learning is well worth further exploration.
@@ IV-B Rethinking and Discussion
Drawing from the aforementioned findings, we argue that the deep models in previous gait recognition methods may not be robust enough. It is caused by the heavy reliance on constrained datasets that struggle with simulating the complexity of outdoor environments. To provide a comprehensive understanding, we elaborate on the following points.
@@
Necessity of Outdoor Evaluation. Previous methods are primarily evaluated on the in-the-lab CASIA-B [10] and OU-MVLP [11]. We argue that this practice suffers from three drawbacks: a) Indoor settings. The walking videos were captured by a camera array, and the subjects were requested to follow a particular route. It makes the data significantly different from that from real-world scenarios. b) Simple background. The simple laboratory background cannot reflect the complex changes of wild scenes. c) Outdated processing methods. The silhouettes were obtained by background subtraction algorithms significantly distinct from the deep-learning-based segmentation methods used for current applications.
Recently, some large-scale real-world gait datasets, such as GREW [8] and Gait3D [9], aim to advance gait recognition from controlled laboratory settings to real scenarios. But designing gait recognition methods based on real-world datasets has still not gained enough attention before the publication of this paper’s conference version.
@@
Figure 4: (a) Overall architecture. The TP and HP respectively denote Temporal and Horizontal Pooling. The Head comprises the separate fully-connected and BNNeck [56] layers, respectively outputting the identity feature vector $f$ and identity distribution vector $p_{\hat{y}}$. (b) Employed blocks, i.e., the 3D and Pseudo 3D residual convolution layers [18, 57, 58]. (c) The proposed three baseline models.
Necessity of Sufficient Ablative Experiments. The ablation study is recognized as the primary means of evaluating the effectiveness of individual components within a specific method. However, we assume that the conclusions drawn from ablative experiments merely on CASIA-B [10] cannot easily extend to real scenarios. The primary reasons are two-fold: a)  CASIA-B contains only 50 subjects for testing. Such few subjects for testing make the results vulnerable to noise. b)  The silhouettes from a good person segmentation model are different from those from the simple background subtraction algorithm of CASIA-B. In principle, performing a comprehensive ablation study across diverse large-scale datasets can yield a robust configuration of hyperparameters.
Necessity of a Strong Backbone. The quality of a model largely hinges on the capability of its backbone. If the backbone is not strong enough, it may inflate the effectiveness of additional modules. The evolution of CNN architectures has progressed from shallow to deep with emerging excellent backbone networks like AlexNet [59], VGG-16 [60], and ResNet [18]. However, previous works [7, 61, 9] in gait recognition have predominantly relied on shallow convolutional neural networks, which are composed of just several convolution layers. As some large-scale real-world datasets [8, 9] become accessible, the necessity of a robust and powerful gait backbone becomes apparent and essential.
V Baselines for Gait Recognition
@@
Building on the insights outlined previously, this section introduces a series of baseline models that reveal several practical principles for in-the-wild gait recognition. Our investigation encompasses appearance-based, model-based, and multi-modal methodologies, contributing to the comprehensiveness of this benchmark study. To ensure broader applicability, we prioritize the development of structurally simple yet experimentally powerful and empirically robust architectures to lay the foundations for follow-ups. To this end, we undertake three-fold efforts:
@@
  • A General Pipeline. To ensure the baseline models’ typicality and universality, a widely recognized pipeline shown in Fig. 4 (a) has been adopted.
@@
  • Two Basic Blocks. To make it simple and scaleable, only two blocks are employed for the network backbone. As illustrated in Fig. 4 (b), they are the 3D residual unit and its pseudo counterpart [57, 58].
@@
  • Three Strong Baselines. As depicted in Fig. 4 (c), the three baselines, DeepGaitV2, SkeletonGait, and SkeletonGait++, share an identical pipeline. They take the silhouette images, skeletal coordinates, and their combination as the input, respectively representing the appearance-based, model-based, and multi-modal baseline model. Notably, we align the data format of silhouettes and skeletons by drawing the latter’s coordinates into an imagery heatmap. It enables a flexible comparison between these two distinct gait modalities. Details are provided in Sec. V-C.
The subsequent content begins by introducing the proposed pipeline, followed by the introductions to the three baselines, DeepGaitV2, SkeletonGait and SkeletonGait++.
V-A The Pipeline